OpenStudy (anonymous):

Ben is making wooden toys for the next arts and crafts sale. Each toy costs Ben $1.80 to make. If he sells the toys for $3.00 each, how many will he have to sell to make a profit of exactly $36.00 ?

OpenStudy (anonymous):

30

OpenStudy (amistre64):

P(x) = 1.20x 36/1.2 = 18/.6 = 180/6 = 30 yep

OpenStudy (anonymous):

sorry, but from where did u get the 1.20x

OpenStudy (anonymous):

3-1.80=1.20

OpenStudy (anonymous):

got it